If 2x + 3 = 5, then value of 3x + 2 is                                        
 If 2x + 3 = 5, then value of 3x + 2 is                                        

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Consider the given equation, 2x + 3 = 5
Transpose 3 to RHS then it becomes -3.

2x = 5 - 3
2x = 2
x = 2/2
x= 1
then value of 3x + 2 is = (3 × 1) + 2
=3 + 2
=5

Thanks